OpenStudy (anonymous):

How many solutions does the equation 4y + 7 = 5 + 2 + 4y have? One Two None Infinitely many

OpenStudy (michaelbp):

one

OpenStudy (anonymous):

thanks

OpenStudy (michaelbp):

medal and fan????

OpenStudy (anonymous):

Correct answer is Infinitely Many. Re-order the equation: 4y - 4y = 5+2-7 -> 4(y-y) = 0. Any value of y will satisfy that equation.